OpenStudy (anonymous):

a barrel contains 101 gallons of water and is being drained at a contant rate of 8 gallons per minutes. write an equation that models the number of gallons.g after t minutes. the equation that models the given situation is____.

OpenStudy (mathstudent55):

In 1 minute, the barrel loses 8 gallons = 8 * 1 gallons In 2 minutes, the barrel loses 16 gallons = 8 * 2 gallons In 3 minutes, the barrel loses 24 gallons = 8 * 3 gallons In t minutes, the barrel loses 8t gallons

OpenStudy (mathstudent55):

The barrel started with 101 gallons, so the equation has to show that 8t gallons are being taken away from 101 gallons. Can you come up with the equation?

OpenStudy (anonymous):

im lost

OpenStudy (mathstudent55):

If you have $101 and $8 is taken away, how would you show that. I don't mean the result. I mean which operation would you use to show $101 take away $8?

OpenStudy (anonymous):

101-8

OpenStudy (anonymous):

93

OpenStudy (anonymous):

ok im lost

OpenStudy (mathstudent55):

Right, $101 take away $8 is $101 - $8 In this case, they want an equation in which you show the full barrel, 101 gallons, and then you take away 8t gallons, so it is: 101 - 8t They also say the number of gallons is represented by the letter g, so the answer is g = 101 - 8t